Find all values of satisfying the inequality \\(-11\\leq 4-3x\\leq 28\\)
1
Expert's answer
2020-06-09T18:35:56-0400

Given "-11 \leq 4-3x \leq 28".

So, "-11 \leq 4-3x \implies -11-4\leq -3x \implies -15 \leq -3x \implies 5 \geq x" .

And, "4-3x \leq 28 \implies -3x\leq 28-4 \implies -3x\leq24 \implies x \geq -8" .

Hence, "-8\leq x\leq 5 \implies x\in[-8,5]."